IR LASER INDUCED CIS ↔ TRANS ISOMERIZATION OF NAPHTHOL IN THE ELECTRONIC EXCITED STATE

K. KOUYAMA, N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578; \underline{T. EBATA}, Department of Chemistry, Graduate School of Science, Hiroshima University, Higashi Hiroshima 739-8526, Japan.